MOLECULAR ION DOPPLER DRIFTS AND ION MOBILITIES IN A GLOW DISCHARGE

\underline{S. CIVIŠ}, and Z. ZELINGER, J. Heyrovský Institute of Physical Chemistry, Academy of Sciences of the Czech Republic, Dolejskova 3, 182 23 Prague 8, Czech Republic.

The mobilities of a number of ions (ArD^{+}, N_{2}H^{+}, SD^{+} and SD^{−}) have been studied on the basis of their Doppler shifts in a positive discharge colum. The technique of infrared velocity modulation and precise microwave measurements ^{a} were employed to investigate the mobilities of spectroscopically identified ions over a broad range of E/N values in an argon buffer gas. For E/N > 100 Td, the reduced mobilities exhibit substantial differencies for particular ions which could not be explained using the classical Langevin and kinetic theories. The experimental results and a simple model based on influence of an inhomogenous electric field will be presented.

aS. Civiš, A. Walters, M. Yu. Tretyakov, S. Bailleux, and M. Bogey, J. Chem. Phys.108, 8369 (1998).